# LingoQA: Video Question Answering for Autonomous Driving

### v2

@@ 5 Empirical Evaluation on LingoQA
With the highly modular architecture of VLMs, the question remains what architectural components of the LingoQA Baseline model and dataset composition contribute the most to its performance? We conduct several ablation studies around the architecture and training paradigm described in Section 4. To this end, we investigate variations to the training strategy, training data composition, frame count, video fusion methods, and the use of different large language models. The results are obtained by having each model generate one answer per question and then compare the predicted answer to the two ground truth answers. Examples of comparisons between our baseline model’s answers and answers from other models from the ablations are presented in Appendix G.
5.1 Training Strategy
@@
The aim of the training strategy experiments is to understand how much the pre-training and the fine-tuning steps contribute to performance. It becomes apparent that fine-tuning leads to improved answers relevant to autonomous driving. The model fine-tuned on the LingoQA dataset has double the performance of the dataset that is pre-trained on generic VQA datasets.
5.2 Training Datasets Mixture
Table 4.2 shows the contributions of our two datasets, Action Dataset and the Scenery Dataset. Both datasets proved influential in improving model performance. We show that fine-tuning on the LingoQA dataset that we open source leads to a considerable improvement compared to general pre-training only.
5.3 Impact of Frame Count
We want to investigate the variation in VQA performance with decreasing and increasing the number of video frames fed into the model. The base model contains 5 frames over a 4-second context. The performance declined when shifting from multi-frame video to a single image representation, which can be explained by the model not getting enough information to answer questions where temporal information is crucial. The performance when increasing the number of frames to 3, 5, and 7 remains relatively consistent. We hypothesize that this is due to the lack of effective video fusion, hence we highlight the opportunity for improved video encoding methods.
5.4 Impact of Video Fusion Strategy
@@
Given how crucial the temporal context is for scenario understanding, this study explores three methods for integrating video frames into the LLM: early-fusion, mid-fusion, and late-fusion. The early-fusion method employs average pooling to condense features from the vision encoder prior to their incorporation into the Q-Former, producing a unified visual feature vector for language space projection. In contrast, the mid-fusion approach, merges video features into fixed-size tokens within the Q-Former with the cross-attention mechanism. The late-fusion method, where individual frame embeddings from Q-Former output are fed into the LLM, allows it to resolve temporal relationships. Our findings demonstrate that both mid-fusion and late-fusion are effective methods for incorporating video content into the model. Mid-fusion allows for a greater number of context tokens through the use of a predetermined number of video embeddings. Conversely, late-fusion shows a slightly enhanced performance by providing comprehensive frame information to the LLM.
5.5 Impact of Large Language Model
@@
We investigate the impact that different Large Language Models have on the overall performance of our vision-language model. As shown in Table 4.2, the best score is achieved by Vicuna-1.5-7B [13] that our base model uses. In the same family of models, Llama-2-7B [48] achieves comparable, but slightly lower performance. Mistral-7B [26], despite its promise of superior performance over Llama-2, proved less effective in our fine-tuning task. Notably, OPT-7B [60] demonstrates significantly lower performance, despite having a similar number of parameters. This discrepancy underscores the crucial role of the pretraining phase in the base language model’s effectiveness.
@@ 5.6 Evaluation of SOTA Vision-Language Models
To demonstrate the relevance of the newly proposed benchmark, we evaluate a series of SOTA vision-language models and compare them to human performance. The models shown in Table 5 are evaluated zero-shot. The best performing models achieve an accuracy as rated by humans of 56.67%, which is below human performance at 93.4%. This highlights the challenging nature of our proposed benchmark.
Table 5: Evaluating vision-language models zero-shot. The performance of existing vision-language models is far from human capability, highlighting the challenging nature of the benchmark.
Human Lingo-J GPT-4 BLEU METEOR CIDEr
Human 93.4 96.6 4.68 81.04 52.92 361.77
GPT-4V 56.67 59.6 3.19 6.30 12.35 42.82
LLaVA 38.97 49.4 2.45 4.23 8.38 38.39
FUYU 17.69 45.4 2.28 1.90 13.00 12.04
6 Discussion
@@
The strength of our contribution comprises proposing a classifier that is highly correlated with human inputs and efficient to run. In conjunction with the evaluation dataset that we propose, it becomes a useful tool for benchmarking vision-language models for autonomous driving on the video question answering task, which has been historically challenging to evaluate in a consistent fashion. With this contribution, autonomous driving research can be accelerated by providing a reliable, efficient, and easy-to-interpret benchmark.
@@
This work still has a few limitations that we discuss and provide guidelines as to how their effects may be mitigated. The classifier has been trained for autonomous driving evaluation and has been shown effective for this purpose. Hence, the classifier can be used for evaluation when paired with the benchmark dataset that we release with the paper. Second, we optimized the classifier to evaluate responses in the style provided by human annotators in the evaluation dataset. The same response style is adopted in the LingoQA training sets and the models. Further details regarding generalisation to response styles is studied in Appendix E. Third, as the classifier is only trained to predict factual correctness, it cannot discern which answer of two equally correct answers humans prefer.
@@ 7 Conclusion
In this paper, we introduced a novel benchmark for Video Question Answering for autonomous driving. The benchmark consists of a evaluation dataset, learned classifier-based metric Lingo-Judge that is highly correlated with human evaluation, a comprehensive high-quality training dataset for autonomous driving. The fast feedback from employing Lingo-Judge facilitates effective exploration in the video QA field. Additionally, the comprehensive experiments on different model combinations presented in this paper can become a foundation for further improvement of end-to-end autonomous driving systems. The LingoQA benchmark is openly released to spur further community research, providing a reliable and highly correlated evaluation method to human ratings.
@@ Acknowledgements
This work was possible through the help of many colleagues across Wayve. We would like to thank Elahe Arani for the valuable discussions and feedback on the report. In particular we would like to acknowledge the support from: Anthony Hu, Miguel Sanchez Lohff, Lorenzo Bertoni, Charlie Lyons-Rothbart, Emma Wang, Harriett-Rose Follas, Kyle Esecson, Ben Foxall, Naama Zahavi, Ruben Diaz, Rudi Rankin, Tilly Pielichaty, Sarah Belghiti, Giulio D’Ippolito, Dan Reisman, Alex Persin, Fergal Cotter, Przemyslaw Mazur, Thomas Sajot, Giacomo Gallino, Alex Garcia Mayans, Tim Geypens, Robin Tweedie, Rebecca Hills.
Appendix A LingoQA Dataset Examples
Further examples on the capabilties existent in the training and the evaluation datasets are shown in Figure 4. The scenery dataset contains highly descriptive elements, such as object colours, junction type, construction zones, traffic lights, and the road layout. The action dataset is complementary and focused on driving competencies, such as the impact of traffic lights on driving and interactions with other road agents. The evaluation dataset contains a broad range of questions aimed to test competencies relevant for autonomous driving. Further examples from the evaluation benchmark are also included in the overview video accompanying the submission. Notably, any personal identifiable information, such as faces and plate ID’s, has been annonymised in the dataset.
Figure 4: LingoQA dataset examples. From left to right: scenery dataset, action dataset, and evaluation dataset.
Appendix B Lingo-Judge Examples
@@
Question
Label
Prediction
CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
How many pedestrians are crossing the road?
Zero pedestrians
There are no pedestrians crossing the road.
23.18 5 5 0.96 True
What is the road speed limit?
20 mph - it is written on the road
The road speed limit is 20 mph.
97.91 4 3 0.95 True
How many cars are driving in your direction?
None
There are no cars driving in my direction.
0.11 4 3 0.96 True
Which vehicle should you follow if any?
The motorcyclist.
If any, I should follow the motorcycle ahead.
0.42 4 5 0.95 True
What is the current action and its justification? Answer in the form “action, justification”
Slow down, there is a stationary van infront of us
I am decelerating because of the stationary truck ahead.
34.79 5 5 0.96 True
What is the current action and its justification? Answer in the form “action, justification”
Stop, Red light
I am stopping because the traffic lights to go straight are red.
18.62 5 5 0.95 True
How many cyclists can you see?
I can see 3 cyclists
I can see two cyclists.
150.29 1 2 0.05 False
What color are the traffic lights showing?
The traffic lights are showing green
The traffic lights are showing red.
329.36 0 1 0.05 False
What action are you taking with respect to the cyclist?
Overtaking them on the right and keeping the speed
I am overtaking the cyclist on the left.
349.52 2 2 0.10 False
In which direction is the bus driving?
The bus is driving in the opposite direction
The bus is driving in the oncoming direction.
404.65 4 5 0.31 False
Are there any parked car on the side of the road?
Yes, there are two cars parked on the right of the road
No, there are no parked cars on either side of the road.
142.40 0 0 0.05 False
Is acceleration necessary in this situation? If so, provide the reason.
No. We should decelerate in this situation because there is a vehicle stopping ahead of us.
No, acceleration is not necessary in this situation as I am already driving at the speed limit.
177.96 3 3 0.31 False
@@
Question
Label
Prediction
CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
Is any vehicle stopped at the zebra crossing?
There is no zebra crossing.
No, there are no vehicles stopped at the zebra crossing.
106.19 4 2 0.93 True
What elements are capturing your attention and what is the reason behind it?
I am making sure there is no vehicle on the left lane I am making my lane change to. I am also looking at the traffic light to make sure it stays green while I am crossing the intersection.
I am paying attention to the bus lane on the left because I need to maintain a safe distance from it.
0.91 3 1 0.57 True
What is the current action and its justification? Answer in the form ”action, justification”
We are taking a left turn, because we are at a t-junction. Then we keep a safe distance from the cyclists in front.
I am decelerating to keep a safe distance from the cyclist ahead of me.
46.45 2 2 0.32 False
@@ Appendix C GPT-4 Grading
Appendix D Lingo-Judge Correlation Study
@@
Appendix E Lingo-Judge Generalisation
To investigate the generalisation abilities of the Lingo-Judge, we examine the performance of the model on a range of answer styles. In particular, we evaluate vision-language models with varying architectures namely GPT4-V, LLaVA and FUYU. Table 8 shows the performance of the Lingo-Judge as measured by the validation accuracy. This highlights that the model performs the best on short, concise answers, akin to those of humans.
Table 8: Robustness to response styles. Investigation into the impact of the response style on validation accuracy.
Lingo-J Val. Acc.
LingoQA 89.50
@@ Appendix E Lingo-Judge Generalisation
few-shot (FS) 83.27
concise (C) 81.63
unprompted (U) 83.06
incorrect (I) 89.59
@@
concise (C) 81.43
unprompted (U) 78.12
FUYU 64.89
@@
Appendix F Training Parameters
@@
Appendix G LingoQA Baseline Examples

### v4

@@ 5.1 Ablation Studies on LingoQA
With the highly modular architecture of VLMs, the question remains what architectural components of the LingoQA Baseline model and dataset composition contribute the most to its performance? We conduct several ablation studies around the architecture and training paradigm described in Section 4. We investigate variations to the training strategy, training data composition, frame count, video fusion methods, and the use of different large language models, as shown in Table 5. The results are obtained by having each model generate one answer per question and then compare the predicted answer to the two ground truth answers. Examples of comparisons between our baseline model’s answers and answers from other models from the ablations are presented in Appendix 0.G.
Training Recipe and Dataset Mixture.
@@
The aim of the training strategy experiments is to understand how much the pre-training and the fine-tuning steps contribute to performance. Fine-tuning on the LingoQA dataset doubles the performance over generic VQA pre-training. The Action Dataset and the Scenery Dataset both prove influential in improving model performance.
Impact of Frame Count.
We investigated the variation in VQA performance with decreasing and increasing the number of video frames fed into the model. The base model contains 5 frames over a 4-second context. The performance declined when shifting from multi-frame video to a single image representation, while remaining close to the multi-frame baseline, showing that a certain proportion of autonomous driving scenarios can be solved from a single frame, as further discussed in Section 5.2. Nonetheless, to fully reach human-level multi-frame performance, video fusion is needed, as shown in Table 5. We conclude that both improved single-frame reasoning and video fusion are required.
Impact of Video Fusion Strategy
@@
This study explores three methods for integrating video frames into the LLM: early-fusion, mid-fusion, and late-fusion. The early-fusion method employs average pooling to condense features from the vision encoder prior to their incorporation into the Q-Former, producing a unified visual feature vector for language space projection. The mid-fusion approach, merges video features into fixed-size tokens within the Q-Former with the cross-attention mechanism. The late-fusion method feeds individual frame embeddings from the Q-Former output into the LLM. Our findings demonstrate that both mid-fusion and late-fusion are effective methods for incorporating video content into the model.
Impact of Large Language Model
@@
We investigate the impact that different Large Language Models have on the overall performance of our vision-language model. The best score is achieved by Vicuna-1.5-7B [13]. In the same family of models, Llama-2-7B [48] achieves comparable, but slightly lower performance. Despite the promise of improved performance, Mistral-7B [26] is less effective in our fine-tuning task. OPT-7B [60] substantially underperforms the others, potentially due to the lower embedding size - 1048 compared to 4096 for all other models.
@@ 5.2 Evaluation of SOTA Vision-Language Models
To demonstrate the relevance of the newly proposed benchmark, we evaluate a series of SOTA vision-language models and compare them to human performance, as shown in Table 5.
Human study.
Human performance is evaluated on both video inputs and a single frame. We find a performance degradation from 96.6% to 81.8% without temporal context. The main failure modes include misclassifying parked cars as engaged in traffic, miscounting developing hazards such as motorcyclists and pedestrians over multiple frames, missing state transitions of traffic lights, and failing to predict the correct speed when behind a vehicle. These results indicate that while video understanding is required for reaching human multi-frame performance, improvements in single-frame reasoning are also crucial.
Fine-tuned models.
Our work identifies a notable 23% performance gap between single-frame LLaVA and single-frame human capability, underlining the benchmark’s significance for the autonomous driving and vision-language community. Our evaluation includes fine-tuning a single-frame LLaVA, a single-frame BLIP-2, and a text-only Vicuna-7B model on LingoQA, as shown in Table 5. Results reveal a performance lower bound of 38% for models lacking visual inputs. Notably, vision-language models, such as BLIP-2 and LLaVA, surpass this text-only baseline by 13% and 20.2%, respectively, with performance enhancements attributed to enhanced perceptual capabilities. Furthermore, LLaVA’s use of a larger CLIP crop size (336) compared to 224 improves performance.
Zero-shot models.
The best zero-shot model GPT-4V performance is still 37% below the multi-frame human performance. This highlights that further advancements are required for frontier models to fully solve the benchmark. The Lingo-Judge accuracy is impacted by the model response style, where long-form incorrect answers may receive high ratings, as it happens with FUYU. Further details are presented in Appendix 0.E.
Table 5: Evaluating vision-language models on LingoQA. The performance of existing vision-language models is far from human capability.
Category No. Frames Human Lingo-J BLEU METEOR CIDEr
Human human study 5 93.3 96.6 81.04 52.92 361.77
Human 1 - 81.8 10.64 15.01 64.45
LingoQA fine-tuned models 5 57.1 60.8 15.00 18.56 65.62
LingoQA 1 - 57.0 14.21 18.40 59.46
LLaVA 1 - 59.0 12.5 18.5 57.0
BLIP-2 1 - 52.2 13.0 17.4 60.1
Vicuna-7B 0 - 38.8 10.1 15.2 51.0
GPT-4V zero-shot models 5 56.61 59.6 6.30 12.35 42.82
LingoQA 5 - 33.6 8.33 14.33 39.16
LLaVA 1 38.97 49.4 4.23 8.38 38.39
FUYU 1 17.69 45.4 1.90 13.00 12.04
6 Discussion and Limitations
@@
The strength of our contribution comprises proposing a classifier that is highly correlated with human inputs and efficient to run. In conjunction with the evaluation dataset that we propose, it becomes a useful tool for benchmarking vision-language models for autonomous driving for visual question answering task, which has been historically challenging to evaluate in a consistent fashion. With this contribution, autonomous driving research can be accelerated by providing a reliable, efficient, and easy-to-interpret benchmark.
@@
The Lingo-Judge is specifically tailored for open-vocabulary evaluation on the LingoQA benchmark. While it demonstrates high speed and accuracy compared to larger models like GPT-4, it is designed akin to the model described in the TruthfulQA [34] paper, which suggests that such specialized models are not expected to generalize well to new questions. Second, we optimized the classifier to evaluate responses in the style provided by human annotators in the evaluation dataset. The same response style is adopted in the LingoQA training sets and the models. Further details regarding generalisation to response styles is studied in Appendix 0.E. Third, as the classifier is only trained to predict factual correctness, it cannot discern which answer of two equally correct answers humans prefer.
@@ 7 Conclusion
In this paper, we introduced a novel benchmark for Visual Question Answering for autonomous driving. The benchmark consists of an evaluation dataset, learned classifier-based metric Lingo-Judge that is highly correlated with human evaluation, a comprehensive high-quality training dataset for autonomous driving. The fast feedback from employing Lingo-Judge facilitates effective exploration in the visual QA field. Additionally, the comprehensive experiments on different model combinations presented in this paper can become a foundation for further improvement of end-to-end autonomous driving systems. The LingoQA benchmark is openly released to spur further community research, providing a reliable and highly correlated evaluation method to human ratings.
@@ Acknowledgements
This work was possible through the help of many colleagues across Wayve. In particular we would like to acknowledge the support from: Anthony Hu, Miguel Sanchez Lohff, Lorenzo Bertoni, Charlie Lyons-Rothbart, Emma Wang, Harriett-Rose Follas, Kyle Esecson, Ben Foxall, Naama Zahavi, Ruben Diaz, Rudi Rankin, Tilly Pielichaty, Sarah Belghiti, Giulio D’Ippolito, Dan Reisman, Alex Persin, Fergal Cotter, Przemyslaw Mazur, Thomas Sajot, Giacomo Gallino, Alex Garcia Mayans, Tim Geypens, Robin Tweedie, Rebecca Hills.
Appendix 0.A LingoQA Dataset Examples
Further examples on the capabilties existent in the training and the evaluation datasets are shown in Figure 4. The scenery dataset contains highly descriptive elements, such as object colours, junction type, construction zones, traffic lights, and the road layout. The action dataset is complementary and focused on driving competencies, such as the impact of traffic lights on driving and interactions with other road agents. The evaluation dataset contains a broad range of questions aimed to test competencies relevant for autonomous driving. Further examples from the evaluation benchmark are also included in the overview video accompanying the submission. The dataset statistics for the evaluation dataset are shown in Figure 5. Notably, any personal identifiable information, such as faces and plate ID’s, has been anonnymised in the dataset.
Figure 4: LingoQA dataset examples. From left to right: scenery dataset, action dataset, and evaluation dataset. Further video examples are provided in the supplementary material accompanying the submission.
Figure 5: Evaluation Dataset Statistics. The evaluation dataset assesses a range of competencies and includes a wide range of objects relevant for autonomous driving.
Appendix 0.B Lingo-Judge Examples
@@
Question Label Prediction CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
How many pedestrians are crossing the road? Zero pedestrians There are no pedestrians crossing the road. 23.18 5 5 0.96 True
What is the road speed limit? 20 mph - it is written on the road The road speed limit is 20 mph. 97.91 4 3 0.95 True
How many cars are driving in your direction? None There are no cars driving in my direction. 0.11 4 3 0.96 True
Which vehicle should you follow if any? The motorcyclist. If any, I should follow the motorcycle ahead. 0.42 4 5 0.95 True
What is the current action and its justification? Answer in the form “action, justification” Slow down, there is a stationary van infront of us I am decelerating because of the stationary truck ahead. 34.79 5 5 0.96 True
What is the current action and its justification? Answer in the form “action, justification” Stop, Red light I am stopping because the traffic lights to go straight are red. 18.62 5 5 0.95 True
How many cyclists can you see? I can see 3 cyclists I can see two cyclists. 150.29 1 2 0.05 False
What color are the traffic lights showing? The traffic lights are showing green The traffic lights are showing red. 329.36 0 1 0.05 False
What action are you taking with respect to the cyclist? Overtaking them on the right and keeping the speed I am overtaking the cyclist on the left. 349.52 2 2 0.10 False
In which direction is the bus driving? The bus is driving in the opposite direction The bus is driving in the oncoming direction. 404.65 4 5 0.31 False
Are there any parked car on the side of the road? Yes, there are two cars parked on the right of the road No, there are no parked cars on either side of the road. 142.40 0 0 0.05 False
Is acceleration necessary in this situation? If so, provide the reason. No. We should decelerate in this situation because there is a vehicle stopping ahead of us. No, acceleration is not necessary in this situation as I am already driving at the speed limit. 177.96 3 3 0.31 False
@@
Question Label Prediction CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
Is any vehicle stopped at the zebra crossing? There is no zebra crossing. No, there are no vehicles stopped at the zebra crossing. 106.19 4 2 0.93 True
What elements are capturing your attention and what is the reason behind it? I am making sure there is no vehicle on the left lane I am making my lane change to. I am also looking at the traffic light to make sure it stays green while I am crossing the intersection. I am paying attention to the bus lane on the left because I need to maintain a safe distance from it. 0.91 3 1 0.57 True
What is the current action and its justification? Answer in the form "action, justification" We are taking a left turn, because we are at a t-junction. Then we keep a safe distance from the cyclists in front. I am decelerating to keep a safe distance from the cyclist ahead of me. 46.45 2 2 0.32 False
@@ Appendix 0.C GPT-4 Grading
Appendix 0.D Lingo-Judge Correlation Study
@@
Appendix 0.E Lingo-Judge Generalisation
To investigate the generalisation abilities of the Lingo-Judge, we examine the performance of the model on a range of answer styles. In particular, we evaluate vision-language models with varying architectures namely GPT4-V, LLaVA and FUYU. We also employ human labelling to obtain ground truth performance of these models. Table 8 shows the performance of the Lingo-Judge as measured by the validation accuracy, as well as human, Lingo-Judge and GPT-4 ratingd for comparison. The Lingo-Judge rates FUYU at 45.4%, compared to 17.69% as rated by humans. However, the Lingo-Judge rates GPT-4V concise at 56.6% compared to 56.67% by humans. This highlights that the model performs the best on short, concise answers, akin to those of humans.
Table 8: Robustness to response styles. Investigation into the impact of the response style on validation accuracy. The Lingo-Judge accuracy is limited on mostly incorrect but long-form answers, with the main failure mode being that these models are rated higher than they should be, compared to the human evaluation.
Human Lingo-Judge GPT-4 Lingo-J Val. Acc.
LingoQA 57.10 60.8 3.30 89.50
@@ Appendix 0.E Lingo-Judge Generalisation
few-shot (FS) 57.69 64.6 3.39 83.27
concise (C) 56.67 59.6 3.19 81.63
unprompted (U) 54.61 59.6 3.24 83.06
incorrect (I) 17.69 26.2 1.55 89.59
@@
concise (C) 38.97 49.4 2.45 81.43
unprompted (U) 28.28 41.8 2.69 78.12
FUYU 17.69 45.4 2.28 64.89
@@
Appendix 0.F Training Parameters
@@
Appendix 0.G LingoQA Baseline Examples